OpenStudy (anonymous):

What is the value of x if a rectangle has a perimeter of 50, a length of x + 5, and width of 11?

OpenStudy (jackellyn):

perimeter is adding all the sides 50=2(x+5)+ 2(11) or 50= (x+5)+(x+5)+11+11 50=2(x+5)+22 Distribute the 2 on the parenthesis 50=2x+10+22 Combine like terms. 50=2x+32 Solve for x, can you tell me what the next step would be to get the x alone?